# Rendezvous in CAVITY: Kinematics and gas properties of an isolated dwarf-dwarf merging pair in a cosmic void region

### v1

@@ 5.1 Main properties of the interacting pair
Our analysis of the gas component, traced by the H emission line in both components, reveals clear signs of rotation in each merging galaxy. This rotation is also evident in their stellar components, suggesting that both merging dwarfs likely possessed rotating discs prior to the merger event. Based on the presence of rotation in each component, along with their high SFR, gas-phase metallicity, and ionized gas, we speculate that these merging pairs were initially star-forming dwarf galaxies with rotating discs. Dwarf galaxies with these characteristics are commonly found in voids (e.g., Kreckel et al., 2011; Domínguez-Gómez et al., 2023).
@@
Analysis of the rotation velocity profiles in Section 4.1.1 indicates that both galaxies have comparable ranges of dynamical mass up to 2.9 kpc from their centres (Fig. 4). Additionally, the nearly identical gas-phase metallicities derived from the integrated spectra of each component, when considered alongside the well-established mass-metallicity relations (e.g., Curti et al., 2020), suggest that the two merging galaxies likely had similar masses. This conclusion is further supported by our stellar mass estimates for each component, based on the fits to the IFU data. Taken together, these findings point to a nearly 1:1 merger event in progress.
@@
The total dynamical mass of this merging pair, based on analysis presented in Fig. 4 is log(Mdyn/M⊙) = 10.22. Durbala et al. (2020) reported a total flux of 1.8 [Jy km s-1] for the HI neutral gas in this system, corresponding to log(MHI/M⊙) = 9.86 0.05, which is 2.2 times smaller than the dynamical mass derived here. We estimated a total stellar mass of log(M⋆/M⊙) 9.7. This value is approximately 1.5 times smaller than the HI mass, making this system possibly one of the rare cases with high abundance of neutral gas studied to date (e.g., Paudel et al., 2018). It is essential to note that uncertainties in the stellar mass estimates and dynamical mass measurements, as discussed in Section 4.2, may impact this comparison. Additionally, the HI profile of this interacting pair, as analysed by Durbala et al. (2020), is incomplete due to missing channels, introducing further uncertainty in the estimated HI mass.
@@
In Section 4.1, we discussed that the gas and stellar components in this merging system are co-rotating in their structures. The individual maps in Fig. 4 suggest that the two components are possibly counter-rotating relative to each other (i.e., component B is retrograde with respect to component A), but confirming this would require knowledge of their 3D orientation on the sky. In retrograde encounters like the one studied here, tidal effects persist for a shorter duration, leading to less extended tidal features. Conversely, major mergers (i.e., mass ratios smaller than 3:1) tend to produce more extended tidal structures that endure longer (e.g., Toomre & Toomre, 1972; Duc & Renaud, 2013). Given that this system is a retrograde merger with a mass ratio of 1:1, it remains unclear which of these factors has played a dominant role in shaping the observed tidal tails in Fig. 1. While resolving this question is beyond the capabilities of our data and the scope of this work, this system serves as an excellent textbook case of a dwarf-dwarf merger, providing an opportunity for modelers to explore the parameter space and assess the impact of internal and orbital dynamics on the formation and evolution of tidal features.
@@
The gas-phase metallicity derived for the entire system and each merging component is lower than what is typically reported for dwarf galaxies of similar stellar masses. This is expected, as recent studies have often found lower gas-phase metallicities, especially in the central regions of galaxies undergoing major mergers or interactions (e.g., Thorp et al., 2019; Garay-Solis et al., 2024). On the other hand, there are a few studies that report higher gas-phase metallicity, by 0.2 dex, in merging/interacting galaxies, particularly in low-mass ones (e.g., Michel-Dansac et al., 2008; Porter et al., 2022). However, the difference between the metallicity measured in this study and values reported for star-forming (non-interacting) dwarf galaxies falls within the systematic uncertainties of approximately 0.2 dex associated with empirical calibrations. As such, no definitive conclusions can be drawn. A systematic analysis of a larger sample of interacting or merging dwarf galaxies, particularly in void environments, is necessary to validate these findings.
We also observed a distinctive distribution of red colours, appearing as two small, connected arcs separating the main body of the galaxies from the faint tidal tail. This red colour distribution (Panel D of Fig. 1), when compared with the dust distribution visible in the DECaLS images (Fig. 1), suggests that it is primarily associated with dust in this interacting system, which is also influenced by the ongoing merger. The dust observed can be related to the ionized gas or the molecular or atomic gas compressed by the merger activity. However, we could not detect this feature in the extinction maps derived from the IFU data, likely due to its lower spatial resolution compared to the INT imaging.
@@
Some studies predict that after coalescence and in the absence of strong environmental influences, merging dwarf galaxies can evolve into more massive, star-forming dwarfs with irregular shapes or even become BCDs (e.g., Bekki, 2008; Kado-Fong et al., 2020; Zhang et al., 2020). Other research suggests that, even after a major merger of progenitor galaxies with gas-rich discs, as seen in CAVITY35843 and CAVITY35844, disc-like morphologies can still emerge in the merger remnant. This outcome is possible only if tidal torquing is less dominant and the gas retains its high angular momentum, allowing it to settle into a disc at the end of the merger (e.g., Springel & Hernquist, 2005; Hopkins et al., 2009). Further investigations into these scenarios for the merging system studied here would require detailed modelling of the merging event, which is beyond the scope of this work.
@@ 5.2 Environment of the interacting system
We excluded these galaxies from the sample and computed the projected number density of the neighbouring galaxies of each interacting dwarf within a distance of 1.5 Mpc (). The distribution of these local number densities is shown in the lower panel of Fig 8. We found that the majority of this sample is located in locally populated regions, with more than three neighbouring galaxies, resembling “group-like” environments. This aligns with previous studies concluding that galaxy mergers are more common in a group-like environment, where galaxies are not as isolated as in voids but also are not moving relatively as fast as those in clusters (e.g., Tran et al., 2008; Alonso et al., 2012; Deason et al., 2014; Paudel et al., 2018; Sureshkumar et al., 2024).
@@
In their study, Paudel et al. (2018) concluded that dwarf-dwarf interactions and mergers occur frequently in low-density environments, based on a search within a 700 kpc projected distance and 700 km s-1 relative velocity difference. While this might appear to contradict the present findings, it actually highlights a bias due to different definitions of local environments for galaxies. As Omori et al. (2023) discussed, the dependence of merger rates on the galaxy environment varies with the scale at which the environment is defined. Our study considered a more extensive velocity difference-projected distance space and lower-mass galaxies as neighbours. Hence, our results indicate that although most galaxies in Paudel et al. (2018) sample do not have massive neighbours (log(M⋆/M⊙) ¿ 10) within a projected distance of 700 kpc, they are not completely isolated.
@@
Based on our definition of the local environment and isolation, only 1% of the Paudel et al. (2018) sample has conditions similar to this merging system, meaning they are nearly 1:1 merging events. Interestingly, most of the isolated merging/interacting cases listed in that study for which the relative mass ratio of merging components (m1/m2) has been reported (10 out of 18 isolated cases) fall in the range of m1/m2 5 making them mostly minor merger events (Michel-Dansac et al., 2008; Paudel et al., 2018). In these cases, the less massive galaxies, usually the dwarfs, merge with the more massive companion, likely due to the larger gravitational forces exerted by the latter. This is not the case for CAVITY35843 and 35844.
@@
The trigger for the merging event between the two dwarf galaxies in this system remains an open question. Most merging events, as discussed before, occur in group-like environments, where high galaxy number density and the tidal field of the host halo promote such interactions. Additionally, isolated mergers typically involve two galaxies with a significant mass difference, where the more massive galaxy traps the less massive one in its gravitational field. However, components A and B in this interacting pair do not follow these patterns. The merging we witness can be due to the global dynamics of the void and the distribution of dark matter within it (e.g., Beygu et al., 2013; Rieder et al., 2013; Courtois et al., 2023) which shape the flow of void galaxies. Chengalur et al. (2017) proposed that the wet mergers in low-density environments, such as voids, take place as a consequence of structure formation, which is more slowly at small scales than in regions with average density. This can also be a possible pathway for the formation of gas-rich discs. Similarly, the past three-body encounters between components of merging system and their nearby dwarf galaxy might be responsible for triggering this merger event. However, the INT deep imaging utilized in this study reveals no tidal streams or features indicative of past three-body encounters. Further studies on the dynamical evolution of voids and mergers involving three-body interactions are needed to evaluate these scenarios.
@@ 6 Summary
In this study, we examined the kinematics and gas properties of CAVITY35843 and CAVITY35844, a dwarf-dwarf merging pair near the centre of a cosmic void. This system is remarkable due to its orientation relative to the line of sight and its unique local and large-scale environment. As an interacting system, it presents a compelling opportunity to explore the interactions typical of gas-rich low-mass galaxies at high redshifts, which are fundamental to the hierarchical galaxy formation model. The main results of this study are summarised below:
@@
While some theoretical studies suggest that major mergers can trigger AGN activity (e.g., Capelo et al., 2017), our analysis using the BPT diagram and the Cid Fernandes et al. (2011) detection method based on EW(H) finds no evidence of AGN activity in either component of this merging pair. However, whether this absence of AGN activity is due to the low-mass nature of the progenitors or the specific merger stage remains uncertain. Some studies suggest that AGN activity peaks in post-coalescence mergers (e.g., Satyapal et al., 2014), but such trends have only been observed in massive merging galaxies (i.e., with log(M/M) ¿ 11.0; Comerford et al., 2024). A more extensive analysis of merging dwarf galaxies at different evolutionary stages is needed to better understand the role of mergers in AGN triggering.
We estimate a neutral gas-to-stellar mass ratio of approximately 1.5, positioning this system among the rare cases exhibiting high neutral gas abundances observed to date (e.g., Paudel et al., 2018; Ellison et al., 2018). While previous studies indicate that dwarf galaxies in low-density environments, such as voids, are generally gas-dominated, our findings further suggest that the neutral gas in this system has persisted through the merging event (at least up to the current stage). High-resolution follow-up investigations of the molecular gas content in this and similar void galaxies are crucial for gaining deeper insight into the complex processes governing mergers.
@@
These results are based on the detailed analysis of a single 1:1 merging dwarf pair located in a void. While this study provides valuable insights, it represents just one case among many that remain to be explored in order to fully grasp the broader properties of dwarf-dwarf mergers across diverse environments. To validate and expand on these findings, future research should target statistically significant samples with well-characterized environmental contexts, utilizing deep IFU spectroscopy and high-quality photometry for more comprehensive analyses.
@@ Acknowledgements.
We thank the referee for their thoughtful comments, which have improved the content and presentation of this paper. This paper is based on data obtained by the CAVITY project, funded by the Spanish Ministry of Science and Innovation under grants PID2020-113689GB-I00 and PID2023-149578NB-I00 as well as by Consejería de Universidad, Investigación e Innovación and Gobierno de España and Unión Europea - NextGenerationEU through grant AST22_4.4, financed by MCIN/AEI/10.13039/501100011033, the project A-FQM-510-UGR20 financed from FEDER/Junta de Andalucía-Consejería de Transformación Económica, Industria, Conocimiento y Universidades/Proyecto and by the grants P20-00334 and FQM108, financed by the Junta de Andalucía (Spain). BB acknowledges financial support from the Grant AST22_4.4, funded by Consejería de Universidad, Investigación e Innovación and Gobierno de España and Unión Europea – NextGenerationEU, and by the research project PID2020-113689GB-I00 and PID2023-149578NB-I00 financed by MCIN/AEI/10.13039/501100011033. SBD acknowledges financial support from the grant AST22.4.4, funded by Consejería de Universidad, Investigación e Innovación and Gobierno de España and Unión Europea – NextGenerationEU, also funded by PID2020-113689GB-I00, financed by MCIN/AEI, and the support of the Spanish Ministry of Science, Innovation and Universities through the project PID–2021–122544NB–C43. IP acknowledges financial support from the grant AST22.4.4, funded by Consejería de Universidad, Investigación e Innovación and Gobierno de España and Unión Europea — NextGenerationEU, and by PID2023-149578NB-I00, financed by MCIN/AEI. AZ, UL, EF, LSM, MAF and MR acknowledge support from projects PID2023-150178NB-I00 and PID2020-114414GB-100, financed by MCIN/AEI/10.13039/501100011033. J.R. acknowledges financial support from the Spanish Ministry of Science and Innovation through the project PID2022-138896NB-C55. M.A-F. acknowledges support from the Emergia program (EMERGIA20_38888) from Consejería de Universidad, Investigación e Innovación de la Junta de Andalucía. DE acknowledges support from a Beatriz Galindo senior fellowship (BG20/00224) from the Spanish Ministry of Science and Innovation. TRL acknowledges support from Juan de la Cierva fellowship (IJC2020-043742-I) and Ramón y Cajal fellowship (RYC2023-043063-I, financed by MCIU/AEI/10.13039/501100011033 and by the FSE+). LSM acknowledges support from Juan de la Cierva fellowship (IJC2019- 041527-I). RGB, acknowledges financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033 and to grant PID2022-141755NB-I00. AFM has received support from RYC2021-031099-I and PID2021-123313NA-I00 of MICIN/AEI/10.13039/501100011033/FEDER,UE, NextGenerationEU/PRT. SDP acknowledges financial support from Juan de la Cierva Formación fellowship (FJC2021-047523-I) financed by MCIN/AEI/10.13039/501100011033 and by the European Union ‘NextGenerationEU’/PRTR, Ministerio de Economía y Competitividad under grants PID2019-107408GB-C44, PID2022-136598NB-C32, and is grateful to the Natural Sciences and Engineering Research Council of Canada, the Fonds de Recherche du Québec, and the Canada Foundation for Innovation for funding. JFB acknowledges support from the PID2022-140869NB-I00 grant from the Spanish Ministry of Science and Innovation. EF acknowledge support from the Junta de Andalucía (Spain) local government through the FQM-108 project. GTR acknowledges financial support from the research project PRE2021-098736, funded by MCIN/AEI/10.13039/501100011033 and FSE+. This research made use of astropy, a community-developed core python (http://www.python.org, Van Rossum & Drake, 2009) package for Astronomy (Astropy Collaboration et al., 2013, 2018, 2022); IPython (Pérez & Granger, 2007); Matplotlib (Hunter, 2007); NumPy (van der Walt et al., 2011); SciPy (Virtanen et al., 2020; Gommers et al., 2024) and Pandas (McKinney, 2010). This research has made use of the NASA/IPAC Extragalactic Database, operated by the Jet Propulsion Laboratory of the California Institute of Technology, under contract with the National Aeronautics and Space Administration. Funding for SDSS-III has been provided by the Alfred P. Sloan Foundation, the Participating Institutions, the National Science Foundation, and the U.S. Department of Energy Office of Science. The SDSS-III web site is http://www.sdss3.org/. The SDSS-IV site is http://www.sdss.org.References

### v2

@@ 5.1 Main properties of the interacting pair
Our analysis of the gas component, traced by the H emission line in both components, reveals clear signs of rotation in each merging galaxy. This rotation is also evident in their stellar components, suggesting that both merging dwarfs likely possessed rotating discs prior to the merger event. Based on the presence of rotation in each component, along with their high SFR, gas-phase metallicity, and ionised gas, we speculate that these merging pairs were initially star-forming dwarf galaxies with rotating discs. Dwarf galaxies with these characteristics are commonly found in voids (e.g. Kreckel et al., 2011; Domínguez-Gómez et al., 2023).
@@
Analysis of the rotation velocity profiles in Section 4.2 indicates that both galaxies have comparable ranges of dynamical mass up to 2.9 kpc from their centres (Fig. 4). Additionally, the nearly identical gas-phase metallicities derived from the integrated spectra of each component, when considered alongside the well-established mass-metallicity relations (e.g. Curti et al., 2020), suggest that the two merging galaxies likely had similar masses. This conclusion is further supported by our stellar mass estimates for each component, based on the fits to the IFU data. Taken together, these findings point to a nearly 1:1 merger event in progress.
@@
The total dynamical mass of this merging pair, based on analysis presented in Fig. 4 is log(Mdyn/M⊙) = 10.22. Durbala et al. (2020) reported a total flux of 1.8 [Jy km s-1] for the HI neutral gas in this system, corresponding to log(MHI/M⊙) = 9.86 0.05, that is 2.2 times smaller than the dynamical mass derived here. We estimated a total stellar mass of log(M⋆/M⊙) 9.7. This value is approximately 1.5 times smaller than the HI mass, making this system possibly one of the rare cases with high abundance of neutral gas studied to date (e.g. Paudel et al., 2018). It is essential to note that uncertainties in the stellar mass estimates and dynamical mass measurements, as discussed in Section 4.3, may impact this comparison. Additionally, the HI profile of this interacting pair, as analysed by Durbala et al. (2020), is incomplete due to missing channels, introducing further uncertainty in the estimated HI mass.
@@
In Section 4.1, we discussed that the gas and stellar components in this merging system are co-rotating in their structures. The individual maps in Fig. 4 suggest that the two components are possibly counter-rotating relative to each other (i.e. component B is retrograde with respect to component A), but confirming this would require knowledge of their 3D orientation on the sky. In retrograde encounters such as the one studied here, tidal effects persist for a shorter duration, leading to less extended tidal features. Conversely, major mergers (i.e. mass ratios smaller than 3:1) tend to produce more extended tidal structures that endure longer (e.g. Toomre & Toomre, 1972; Duc & Renaud, 2013). Given that this system is a retrograde merger with a mass ratio of 1:1, it remains unclear which of these factors has played a dominant role in shaping the observed tidal tails in Fig. 1. While resolving this question is beyond the capabilities of our data and the scope of this work, this system serves as an excellent textbook case of a dwarf-dwarf merger, providing an opportunity for modellers to explore the parameter space and assess the impact of internal and orbital dynamics on the formation and evolution of tidal features.
@@
The gas-phase metallicity derived for the entire system and each merging component is lower than what is typically reported for dwarf galaxies of similar stellar masses. This is expected, as recent studies have often found lower gas-phase metallicities, especially in the central regions of galaxies undergoing major mergers or interactions (e.g. Thorp et al., 2019; Garay-Solis et al., 2024). On the other hand, there are a few studies that report higher gas-phase metallicity, by 0.2 dex, in merging and interacting galaxies, particularly in low-mass ones (e.g. Michel-Dansac et al., 2008; Porter et al., 2022). However, the difference between the metallicity measured in this study and values reported for star-forming (non-interacting) dwarf galaxies falls within the systematic uncertainties of approximately 0.2 dex associated with empirical calibrations. As such, no definitive conclusions can be drawn. A systematic analysis of a larger sample of interacting or merging dwarf galaxies, particularly in void environments, is necessary to validate these findings.
We also observed a distinctive distribution of red colours, appearing as two small, connected arcs separating the main body of the galaxies from the faint tidal tail. This red colour distribution (Panel D of Fig. 1), when compared with the dust distribution visible in the DECaLS images (Fig. 1), suggests that it is primarily associated with dust in this interacting system, that is also influenced by the ongoing merger. The dust observed can be related to the ionised gas or the molecular or atomic gas compressed by the merger activity. However, we could not detect this feature in the extinction maps derived from the IFU data, likely due to its lower spatial resolution compared to the INT imaging.
@@
Some studies predict that after coalescence and in the absence of strong environmental influences, merging dwarf galaxies can evolve into more massive, star-forming dwarfs with irregular shapes or even become BCDs (e.g. Bekki, 2008; Kado-Fong et al., 2020; Zhang et al., 2020). Other research suggests that, even after a major merger of progenitor galaxies with gas-rich discs, as seen in CAVITY35843 and CAVITY35844, disc-like morphologies can still emerge in the merger remnant. This outcome is possible only if tidal torquing is less dominant and the gas retains its high angular momentum, allowing it to settle into a disc at the end of the merger (e.g. Springel & Hernquist, 2005; Hopkins et al., 2009). Further investigations into these scenarios for the merging system studied here would require detailed modelling of the merging event, which is beyond the scope of this work.
@@ 5.2 Environment of the interacting system
We excluded these galaxies from the sample and computed the projected number density of the neighbouring galaxies of each interacting dwarf within a distance of 1.5 Mpc (). The distribution of these local number densities is shown in the lower panel of Fig 8. We found that the majority of this sample is located in locally populated regions, with more than three neighbouring galaxies, resembling ‘group-like’ environments. This aligns with previous studies concluding that galaxy mergers are more common in a group-like environment, where galaxies are not as isolated as in voids but also are not moving relatively as fast as those in clusters (e.g. Tran et al., 2008; Alonso et al., 2012; Deason et al., 2014; Paudel et al., 2018; Sureshkumar et al., 2024).
@@
In their study, Paudel et al. (2018) concluded that dwarf-dwarf interactions and mergers occur frequently in low-density environments, based on a search within a 700 kpc projected distance and 700 km s-1 relative velocity difference. While this might appear to contradict the present findings, it actually highlights a bias due to different definitions of local environments for galaxies. As Omori et al. (2023) discussed, the dependence of merger rates on the galaxy environment varies with the scale at which the environment is defined. Our study considered a more extensive velocity difference-projected distance space and lower-mass galaxies as neighbours. Hence, our results indicate that although most galaxies in the Paudel et al. (2018) sample do not have massive neighbours (log(M⋆/M⊙) ¿ 10) within a projected distance of 700 kpc, they are not completely isolated.
@@
Based on our definition of the local environment and isolation, only 1% of the Paudel et al. (2018) sample has conditions similar to this merging system, meaning they are nearly 1:1 merging events. Interestingly, most of the isolated merging and interacting cases listed in that study for which the relative mass ratio of merging components (m1/m2) has been reported (10 out of 18 isolated cases) fall in the range of m1/m2 5 making them mostly minor merger events (Michel-Dansac et al., 2008; Paudel et al., 2018). In these cases, the less massive galaxies, usually the dwarfs, merge with the more massive companion, likely due to the larger gravitational forces exerted by the latter. This is not the case for CAVITY35843 and 35844.
@@
The trigger for the merging event between the two dwarf galaxies in this system remains an open question. Most merging events, as discussed before, occur in group-like environments, where high galaxy number density and the tidal field of the host halo promote such interactions. Additionally, isolated mergers typically involve two galaxies with a significant mass difference, where the more massive galaxy traps the less massive one in its gravitational field. However, components A and B in this interacting pair do not follow these patterns. The merging we witness can be due to the global dynamics of the void and the distribution of dark matter within it (e.g. Beygu et al., 2013; Rieder et al., 2013; Courtois et al., 2023) which shape the flow of void galaxies. Chengalur et al. (2017) proposed that the wet mergers in low-density environments, such as voids, take place as a consequence of structure formation, which is more slowly at small scales than in regions with average density. This can also be a possible pathway for the formation of gas-rich discs. Similarly, the past three-body encounters between components of merging system and their nearby dwarf galaxy might be responsible for triggering this merger event. However, the INT deep imaging utilised in this study reveals no tidal streams or features indicative of past three-body encounters. Further studies on the dynamical evolution of voids and mergers involving three-body interactions are needed to evaluate these scenarios.
@@ 6 Summary
In this study, we examined the kinematics and gas properties of CAVITY35843 and CAVITY35844, a dwarf-dwarf merging pair near the centre of a cosmic void. This system is remarkable due to its orientation relative to the line of sight and its unique local and large-scale environment. As an interacting system, it presents a compelling opportunity to explore the interactions typical of gas-rich low-mass galaxies at high redshifts, that are fundamental to the hierarchical galaxy formation model. The main results of this study are summarised below:
@@
While some theoretical studies suggest that major mergers can trigger AGN activity (e.g. Capelo et al., 2017), our analysis using the BPT diagram and the Cid Fernandes et al. (2011) detection method based on EW(H) finds no evidence of AGN activity in either component of this merging pair. However, whether this absence of AGN activity is due to the low-mass nature of the progenitors or the specific merger stage remains uncertain. Some studies suggest that AGN activity peaks in post-coalescence mergers (e.g. Satyapal et al., 2014), but such trends have only been observed in massive merging galaxies (i.e. with log(M⋆/M⊙) ¿ 11.0; Comerford et al., 2024). A more extensive analysis of merging dwarf galaxies at different evolutionary stages is needed to better understand the role of mergers in AGN triggering.
We estimate a neutral gas-to-stellar mass ratio of approximately 1.5, positioning this system among the rare cases exhibiting high neutral gas abundances observed to date (e.g. Paudel et al., 2018; Ellison et al., 2018). While previous studies indicate that dwarf galaxies in low-density environments, such as voids, are generally gas-dominated, our findings further suggest that the neutral gas in this system has persisted through the merging event (at least up to the current stage). High-resolution follow-up investigations of the molecular gas content in this and similar void galaxies are crucial for gaining deeper insight into the complex processes governing mergers.
@@
These results are based on the detailed analysis of a single 1:1 merging dwarf pair located in a void. While this study provides valuable insights, it represents just one case among many that remain to be explored in order to fully grasp the broader properties of dwarf-dwarf mergers across diverse environments. To validate and expand on these findings, future research should target statistically significant samples with well-characterised environmental contexts, utilising deep IFU spectroscopy and high-quality photometry for more comprehensive analyses.
@@ Acknowledgements.
We thank the referee for their thoughtful comments, which have improved the content and presentation of this paper. This paper is based on data obtained by the CAVITY project, funded by the Spanish Ministry of Science and Innovation under grants PID2020-113689GB-I00 and PID2023-149578NB-I00 as well as by Consejería de Universidad, Investigación e Innovación and Gobierno de España and Unión Europea - NextGenerationEU through grant AST22_4.4, financed by MCIN/AEI/10.13039/501100011033, the project A-FQM-510-UGR20 financed from FEDER/Junta de Andalucía-Consejería de Transformación Económica, Industria, Conocimiento y Universidades/Proyecto and by the grants P20-00334 and FQM108, financed by the Junta de Andalucía (Spain). Based on observations collected at Centro Astronómico Hispano en Andalucía (CAHA) at Calar Alto, proposal F21-3.5-101, operated jointly by Junta de Andalucía and Consejo Superior de Investigaciones Científicas (IAA-CSIC). BB acknowledges financial support from the Grant AST22_4.4, funded by Consejería de Universidad, Investigación e Innovación and Gobierno de España and Unión Europea – NextGenerationEU, and by the research project PID2020-113689GB-I00 and PID2023-149578NB-I00 financed by MCIN/AEI/10.13039/501100011033. SBD acknowledges financial support from the grant AST22.4.4, funded by Consejería de Universidad, Investigación e Innovación and Gobierno de España and Unión Europea – NextGenerationEU, also funded by PID2020-113689GB-I00, financed by MCIN/AEI, and the support of the Spanish Ministry of Science, Innovation and Universities through the project PID–2021–122544NB–C43. IP acknowledges financial support from the grant AST22.4.4, funded by Consejería de Universidad, Investigación e Innovación and Gobierno de España and Unión Europea — NextGenerationEU, and by PID2023-149578NB-I00, financed by MCIN/AEI. AZ, UL, EF, LSM, MAF and MR acknowledge support from projects PID2023-150178NB-I00 and PID2020-114414GB-100, financed by MCIN/AEI/10.13039/501100011033. J.R. acknowledges financial support from the Spanish Ministry of Science and Innovation through the project PID2022-138896NB-C55. M.A-F. acknowledges support from the Emergia program (EMERGIA20_38888) from Consejería de Universidad, Investigación e Innovación de la Junta de Andalucía. DE acknowledges support from a Beatriz Galindo senior fellowship (BG20/00224) from the Spanish Ministry of Science and Innovation. TRL acknowledges support from Juan de la Cierva fellowship (IJC2020-043742-I) and Ramón y Cajal fellowship (RYC2023-043063-I, financed by MCIU/AEI/10.13039/501100011033 and by the FSE+). LSM acknowledges support from Juan de la Cierva fellowship (IJC2019- 041527-I). RGB, acknowledges financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033 and to grant PID2022-141755NB-I00. AFM has received support from RYC2021-031099-I and PID2021-123313NA-I00 of MICIN/AEI/10.13039/501100011033/FEDER,UE, NextGenerationEU/PRT. SDP acknowledges financial support from Juan de la Cierva Formación fellowship (FJC2021-047523-I) financed by MCIN/AEI/10.13039/501100011033 and by the European Union ‘NextGenerationEU’/PRTR, Ministerio de Economía y Competitividad under grants PID2019-107408GB-C44, PID2022-136598NB-C32, and is grateful to the Natural Sciences and Engineering Research Council of Canada, the Fonds de Recherche du Québec, and the Canada Foundation for Innovation for funding. JFB acknowledges support from the PID2022-140869NB-I00 grant from the Spanish Ministry of Science and Innovation. EF acknowledge support from the Junta de Andalucía (Spain) local government through the FQM-108 project. GTR acknowledges financial support from the research project PRE2021-098736, funded by MCIN/AEI/10.13039/501100011033 and FSE+. This research made use of astropy, a community-developed core python (http://www.python.org, Van Rossum & Drake, 2009) package for Astronomy (Astropy Collaboration et al., 2013, 2018, 2022); IPython (Pérez & Granger, 2007); Matplotlib (Hunter, 2007); NumPy (van der Walt et al., 2011); SciPy (Virtanen et al., 2020; Gommers et al., 2024) and Pandas (McKinney, 2010). This research has made use of the NASA/IPAC Extragalactic Database, operated by the Jet Propulsion Laboratory of the California Institute of Technology, under contract with the National Aeronautics and Space Administration. Funding for SDSS-III has been provided by the Alfred P. Sloan Foundation, the Participating Institutions, the National Science Foundation, and the U.S. Department of Energy Office of Science. The SDSS-III web site is http://www.sdss3.org/. The SDSS-IV site is http://www.sdss.org.References